OpenStudy (anonymous):

Use identities (no calculators) to fi nd the exact value for sin (pi/12)cos (pi/12) + cos (pi/12)sin(pi/12)

OpenStudy (anonymous):

again? this is \[sin(a+b)\] for \[a=b=\frac{\pi}{12}\]

OpenStudy (anonymous):

\[\frac{\pi}{12}+\frac{\pi}{12}=\frac{\pi}{6}\]

OpenStudy (anonymous):

\[sin(\frac{\pi}{6})=\frac{1}{2}\]